# Considering momentum dependence of heavy quarks interaction rate in calculating drag coefficient, and its impact on energy loss of b and c quarks

### v1

@@ Abstract
This paper examines the effect of heavy particle momentum on their scattering frequency and calculates the drag coefficient by expanding the energy loss terms as polynomial expansions of momentum ($p$). This approach allows for a more precise investigation of momentum dependence of drag coefficient by incorporating the linear terms of these expansions. Additionally, the influence of particle’s momentum on radiative and collisional energy loss is more clearly determined. The study focuses on calculation of the nuclear modification factor ($R_{AA}$) of charm and bottom quarks in Pb-Pb collisions at $\sqrt{S_{NN}}=5.02\>TeV$. The results are compared with the latest data from ALICE and ATLAS experiments, conducted in 2021 and 2022. In this article, the initial distribution functions have been evolved numerically based on the Fokker-Planck equation.
@@ I Introduction
The main goal of ultra-relativistic heavy ion collisions at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) is to investigate QCD matter under extreme conditions. These experiments involve the collision of heavy ions, such as gold or lead, at very high energies to create a hot, dense environment of deconfined quarks and gluons, known as the Quark Gluon Plasma (QGP) STAR:2005gfr ; PHENIX:2004vcz ; QGP.Collins:1974ky . This highly excited state of matter, whose main constituents are light quarks and gluons, exhibits properties similar to a nearly perfect fluid. Therefore, it can be effectively described using hydrodynamic models while the system reaches local thermal equilibrium Ollitrault:2007du ; Becattini:2014rea .
@@
Heavy quarks (such as charm and bottom) play an essential role in studying the properties of quark-gluon plasma as they are best probe for determining the medium’s transport properties HQ.vanHees:2005wb ; HQ.Rapp:2009my . These quarks are too heavy to be significantly produced through the interaction of thermal particles in the QGP. Therefore, the dominant mechanism for their production arises from hard interactions between partons within the colliding nuclei. Moreover, due to their large mass, these particles require a longer duration to reach thermal equilibrium with the environment. In some cases, they may even leave the plasma before achieving equilibrium. Thus, they are valuable probes to study the entire spacetime history of the deconfined medium. To gain a deeper insight into the properties of heavy quarks re-scattering in the QGP, one can analyze the evolution of their distribution function over time in the transverse momentum ($P_{T}$) plane using the Fokker-Planck (FP) equation.
Heavy quarks experience energy loss as they traverse within the hot, dense medium created by collisions. This energy dissipation occurs through two mechanisms: elastic collisions with constituents of the QGP and gluon bremsstrahlung or radiation resulting from interaction with other quarks, antiquarks, and gluons in the thermal environment. The interaction between heavy quarks and dynamic thermal media is characterized by transport coefficients. The drag coefficient, which quantifies the resistance experienced by a parton while moving through the plasma, is directly related to energy loss and can be calculated from the rate of energy loss per unit distance. Considerable efforts have been dedicated to achieving more accurate calculations of the drag and diffusion coefficients. One aspect of these researches examine the momentum dependence of the drag coefficient Mazumder:2011nj and momentum or temperature dependence of diffusion coefficient Hong:2023cwl ; Li:2019lex . Another part of these efforts is concentrated on developing various models for energy dissipation. There are several formulations available for calculating collisional Peshier:2006hi ; Braaten:1991jj ; Peigne:2008nd ; J. D. Bjorken ; Braaten:1991we and radiative energy loss Armesto:2003jh ; Armesto:2004vz ; Gyulassy:2000fs ; Djordjevic:2003zk . For a comprehensive review of many of these formalisms see Ref. Jamil:2010te .
To accurately calculate the drag coefficient, it is essential to consider the number of interactions that heavy particles undergo during propagation and the amount of energy lost in each interaction. As particles propagate through the thermal bath, their energy and momentum decrease, leading to a reduction in the number of both elastic and inelastic collisions over time. In this article, we determine the drag coefficient by considering energy loss coefficients as a linear function of particle’s momentum. This approach examines the impact of particle’s momentum on its collision rate, allowing for a more precise analysis of collisional and radiative energy losses. In this way, we analyze the drag coefficient as a function of momentum, addressing limitations in previous studies that overlooked or simplified its momentum dependence and lost some details.
@@
This paper is organized as follows: In the next section, we provide a brief overview of the Fokker-Planck equation. Additionally, we discuss the space-time evolution of the system and outline the initial conditions. In Section (III), the drag and diffusion coefficients for collisional and radiative processes have been discussed. Our results are presented in Section (IV), and section (V) is dedicated to summary and discussions.
@@ II System evolution
As mentioned before, heavy quarks reach thermodynamic equilibrium more slowly compared to lighter particles. Consequently, the interaction between heavy quarks and medium can be considered as interaction between equilibrium and non-equilibrium states. The equilibrium component, known as Quark-Gluon Plasma, is presumed to form at an initial temperature ($T_{i}$) and time ($\tau_{i}$) after the nuclear collision. The QGP expands and cools due to its high internal pressure, until transition to the hadronic phase at temperature $T_{c}$. Heavy quarks as the non-equilibrium component, exhibit Brownian motion within the thermal environment of the QGP Moore:2004tg ; Alam:1994sc . Therefore, Fokker-Planck equation, which is a simplified version of the Boltzmann equation for small momentum transformations Svetitsky:1987gq ; Rapp:2009my , provides a suitable framework for investigating the evolution of the heavy quark’s momentum distribution. According to this equation we have:
$\frac{\partial}{\partial t}f(p,t)=-\frac{\partial}{\partial p_{i}}\left[A_{i}(p)f(p,t)\right]+\frac{\partial^{2}}{\partial p_{i}\partial p_{j}}\left[D_{ij}(p)f(p,t)\right].$ (1)
Where $f(p,t)$ stands for the distribution function of heavy quarks and $A_{i}(p)$ and $D_{ij}(p)$ are related to the drag and diffusion coefficients, respectively. These coefficients characterize the interaction of heavy quarks with the QGP and are generally influenced by both temperature and momentum. The i and j indices represent different spatial directions, which can be omitted as we are considering a spatially uniform quark-gluon plasma in transverse plane. In the next section, we will discuss these coefficients in detail.
@@
[…]
@@
[…]
@@ III Drag and diffusion coefficients
To solve the Fokker-Planck equation and derive the distribution function of heavy quarks at the transition temperature ($T_{c}$), we require the drag and diffusion coefficients. The diffusion coefficient can be calculated utilizing Einstein’s relation under conditions of weak coupling between the heavy quarks and the thermal bath Diffusion.Srivastava:2016igg ; Diffusion.Das:2010tj ; Diffusion.Akamatsu:2008ge . According to this relation, we have:
$D(p)=TA(p)E$ (4)
T is the temperature of the thermal bath and E is the energy of heavy quarks. Therefore, by calculating the drag coefficient, we will have both the drag and diffusion coefficients.
It is important to note that the drag coefficient carries information about the dynamics of heavy quark interactions with the medium and is influenced by the properties of the thermal bath. Consequently, the key aspect of determining the temporal evolution of heavy quark distribution function lies in calculating the drag force acting on them, or equivalently, the rate of energy loss per unit distance traveled by the heavy quark in the medium. So, we have:
@@
We have considered both kind of energy loss in calculating the drag coefficient. Therefore, it can be determined through the following relation:
@@
$k_{1}$ and $k_{2}$ are typically treated as constants, which can be derived from the best fit to the experimental results. These constants account for the number of collisions, degeneracy, and other factors.
However, to precisely determine the drag coefficient, it is necessary to consider the impact of particles’ momentum on the rate of their collisions, which results in energy loss. As the kinetic energy and momentum of heavy quarks decrease during propagation in QGP, the frequency of their collisions with plasma constituents is expected to decrease as well. Therefore, it can be said that the number of elastic and inelastic interactions, depends on the momentum of the particles.
Consequently, in this article, we have formulated the drag coefficient as follows:
@@
The functions $k^{\prime}(p)$ and $k^{\prime\prime}(p)$ represent collisional and radiative energy loss coefficients, respectively. We assume these functions comprised of a constant value and a modified value which is expressed as polynomial expansions of momentum, as follows:
[…]
We consider these expansions up to first-order which include linear terms. By considering the energy loss coefficients as a function of $p$, the impact of particle’s momentum on the number of its interactions leading to collisional (elastic) and radiative (inelastic) energy losses can be examined separately. Setting $\alpha$ and $\beta$ to zero yields the initial expression of the drag coefficient in the equation (6).
@@
All coefficients including $k_{1}$, $k_{2}$, $\alpha$ and $\beta$ are determined through the fitting of experimental data. Figure (1) presents the drag coefficient for both cases of momentum-dependent coefficients and constant coefficients, at $T=250\;MeV$. As shown in the figure, incorporating the second expansion terms in the drag coefficient calculation results in a significant increase in its value at higher momenta. Consequently, we anticipate greater energy suppression in these regions. By comparing the $R_{AA}$ results obtained from above equation with the case where the energy loss coefficients are considered as constant values (expanded to the first term), the degrees of variation or improvement in the results can be measured.
In order to calculate collisional and radiative energy loss, a variety of methods have been introduced in recent studies. In our previous work Nezhad:2023mvb , we evaluated and compared three different approaches for calculating collisional energy loss, and found that the approach proposed by Braaten and Thoma, which utilizes the Hard-Thermal-Loop (HTL) framework Braaten:1991we , aligns better with recent experimental data. Therefore, we have employed the same method for calculating collisional energy loss in this work. Radiative energy loss has been calculated using the reaction operator formalism (DGLV) Gyulassy:2000fs ; Djordjevic:2003zk ; Wicks:2005gt and the generalized dead cone approach Saraswat:2015yll ; Saraswat:2017vuy .
Figure 1: Drag coefficient with momentum-dependent coefficients and constant coefficients at T=250 MeV.
@@ IV Results
The nuclear suppression factor, $R_{AA}$, defined as the ratio of quenched to unquenched $P_{T}$ distribution, is derived from experimental measurements. This quenching phenomenon results from the interaction between heavy quarks and medium. This factor is influenced by several key processes, including the initial production of heavy quarks, the evolution of the medium, hadronization, and the interactions of heavy quarks within quark-gluon plasmas. We concentrate on the momentum-dependent drag coefficient and energy loss effects in QGP and estimate $R_{AA}$ as the ratio of the final heavy quark distribution in heavy-ion collisions to the initial heavy quark distribution in proton-proton collisions:
@@
The proportionality sign can be replaced with an equal sign by considering the ratio of the number of collisions and also degrees of degeneracy. We will adjust this coefficient by optimizing procedure. In this section, we calculate the nuclear suppression factor of charm and bottom quarks in a Pb-Pb collision at a center-of-mass energy of 5.02 TeV and compare our findings with recent LHC experimental data collected in 2021 and 2022. The parton distribution functions are evaluated using the Fokker-Planck equation, which we solve numerically at third-order relativistic hydrodynamics until the fireball reaches it’s freeze-out temperature of $T_{c}=155$ MeV. We calculate drag coefficient for both cases of first-order momentum expansion (9) and as a constant term (6). This procedure allows us to compare the results with experimental data and investigate the impact of momentum dependence of energy loss coefficients on results.
The final results are optimized to align with experimental data by adjusting free parameters like $\alpha$ and $\beta$ and minimizing the unweighted Chi-squared value:
@@
$R_{AA}^{exp}$ and $R_{AA}^{th}$ represent experimental and theoretical predictions for the suppression factor, respectively, while $\sigma$ indicates experimental error.
In our parameter optimization process, we utilize the Minuit package Minuit.James:1975dr , which is an effective and powerful tool that enables us to fit more free parameters and achieve high accuracy in minimizing the chi-squared value. This resource is essential for ensuring that our optimization efforts are both precise and reliable, thus enhancing the overall quality of our results.
Figure 2: Nuclear modification factor of charm quark in the (20–40%) centrality Pb-Pb collisions at $\sqrt{S_{NN}}=5.02$ TeV compared with ALICE 2021 data at forward rapidity ($42.5<y<$ 4) ALICE:2020sjb . The solid line represents the results obtained from the momentum-dependent coefficients, and dashed line corresponds to the drag coefficient with constant coefficients.
@@
Figures (2) and (3) illustrate $R_{AA}$ results for charm quark, derived from fitting theoretical predictions to the ALICE 2021 ALICE:2020sjb and ALICE 2022 data ALICE:2021rxa , respectively. Analysis of both figures indicates that the plot derived from the momentum-dependent drag coefficient, exhibits a more precise alignment with the experimental data. Furthermore, the $\chi^{2}$ value decreases from 0.3 to 0.2 for ALICE 2022 dataset and decreases from 0.47 to 0.44 for ALICE 2021 dataset, when we use momentum dependence coefficients.
@@
The coefficients $\alpha$ and $\beta$ in Equation (9) for the c quark are found to be $\alpha=0.02$ and $\beta=0.05$. These values indicate that at higher momentum, both collisional and radiative energy losses of particles increase. Specifically, the impact of particle’s momentum on radiative energy loss becomes more significant and should be encountered. We also obtain $k_{1}=1.5$ and $k_{2}=2$ through the optimization process, which implies that the radiative energy loss for charm quarks is consistently greater.
Figure (4) shows our result for b quark fitted on ATLAS 2022 data ATLAS:2021xtw . Due to the limited amount of experimental data for the b quark, it is not feasible to accurately fit a model with too many free parameters. Therefore, in calculating drag coefficient for b quark, we have only considered the first term of the expansion with constant coefficients. By comparing theoretical results with the experimental data, the constant coefficients in (6) are found to be $k_{1}=5.5$ and $k_{2}=1.8$ Therefore, it seems that the collisional energy loss for b quark, is significantly higher than the radiative energy loss. We also have $\chi^{2}=0.124$ and as illustrated in the figure, the theoretical results exhibit a very good agreement with the experimental data.
As previously mentioned, examining higher-order terms in the drag coefficient is more challenging for b quark. However, by simulation and generating additional data point, the parameters $\alpha$ and $\beta$ for this quark are found to be similar to those of c quark. The difference is that for the b quark, even by considering higher order coefficients, we find $k_{1}$ greater than $k_{2}$. The values of $k_{1}$ are found to lie within the approximate range of $32.5<k_{1}<$ 3 and $k_{2}$ is in the approximate range of $1.81<k_{2}<$ 1.8 . These values indicates that for the bottom quark, collisional energy loss is dominant mechanism for both calculation of the drag coefficient. This is mainly due to its large mass which suppresses gluon radiation Dokshitzer:2001zm .
@@
Figure 4: Nuclear modification factor of bottom quark in the (40–60%) centrality Pb-Pb collisions at $\sqrt{S_{NN}}=5.02$ TeV compared with ATLAS 2022 data ATLAS:2021xtw . This graph is obtained from considering the drag coefficient with constant coefficients.

### v3

# Impact of momentum-dependent drag coefficient on energy loss of charm and bottom quarks in QGP

@@ Abstract
This paper investigates the influence of heavy-particle momentum on their interaction rate and the resulting drag coefficient in a quark-gluon plasma. To account for this momentum dependence, the drag coefficient is derived by expressing the energy loss coefficients as polynomial expansions of momentum. This approach allows for a more precise investigation of the momentum dependence of the drag coefficient by incorporating linear terms into the expansion. Furthermore, the effects of particle momentum on radiative and collisional energy loss are determined more accurately. The study focuses on calculating the nuclear modification factor ($R_{AA}$) of charm and bottom quarks in Pb-Pb collisions at $\sqrt{S_{NN}}=5.02\>TeV$. The initial distribution functions are evolved numerically using the Fokker–Planck equation. The results are compared with the latest experimental data from ALICE and ATLAS, collected in 2021 and 2022.
@@ I Introduction
The main goal of ultra-relativistic heavy ion collisions at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) is to investigate QCD matter under extreme conditions. These experiments involve the collision of heavy ions, such as gold or lead, at very high energies to create a hot, dense environment of deconfined quarks and gluons, known as the Quark-Gluon Plasma (QGP) STAR:2005gfr ; PHENIX:2004vcz ; QGP.Collins:1974ky . This highly excited state of matter, whose main constituents are light quarks and gluons, exhibits properties similar to a nearly perfect fluid. Therefore, it can be effectively described using hydrodynamic models while the system reaches local thermal equilibrium Ollitrault:2007du ; Becattini:2014rea .
@@
Heavy quarks (such as charm and bottom) play an essential role in studying the properties of a quark-gluon plasma, as they are the best probe for determining the medium’s transport properties HQ.vanHees:2005wb ; HQ.Rapp:2009my . These quarks are too heavy to be significantly produced through the interaction of thermal particles in the QGP. Therefore, the dominant mechanism for their production arises from hard interactions between partons within the colliding nuclei. Moreover, due to their large mass, these particles require a longer duration to reach thermal equilibrium with the environment. In some cases, they may even leave the plasma before achieving equilibrium. Thus, they are valuable probes to study the entire spacetime history of the deconfined medium. To gain a deeper insight into the properties of heavy quarks re-scattering in the QGP, one can analyze the evolution of their distribution function over time in the transverse momentum ($P_{T}$) plane.
Heavy quarks experience energy loss as they propagate through the hot, dense medium created by collisions. This energy dissipation occurs through two primary mechanisms: elastic scatterings with the constituents of the QGP and gluon bremsstrahlung or radiation resulting from interaction with thermal quarks and gluons. From a transport-theoretical perspective, the interaction between heavy quarks and the dynamically evolving QGP can be characterized by transport coefficients, which encode the cumulative effect of multiple microscopic interactions with the medium. Among these coefficients, the drag coefficient plays a central role, as it quantifies the average resistance experienced by a parton while traversing the plasma and is directly related to its energy loss per unit path length.
Extensive theoretical efforts have been devoted to obtaining more accurate estimates of the drag and diffusion coefficients. In particular, several studies have investigated the momentum dependence of the drag coefficient Mazumder:2011nj ; Prakash:2023hfj , as well as the momentum or temperature dependence of the diffusion coefficient Hong:2023cwl ; Li:2019lex . In parallel, significant progress has been made in developing theoretical models for heavy-quark energy dissipation in the medium, including both collisional Peshier:2006hi ; Braaten:1991jj ; Peigne:2008nd ; J. D. Bjorken ; Braaten:1991we and radiative Armesto:2003jh ; Armesto:2004vz ; Gyulassy:2000fs ; Djordjevic:2003zk energy loss mechanisms. For a comprehensive review of many of these formalisms, see Ref. Jamil:2010te .
An accurate determination of the drag coefficient requires accounting for both the frequency of interactions experienced by heavy quarks during their propagation and the average amount of energy lost in each interaction. As heavy quarks propagate through the thermal bath, their energy and momentum decrease, leading to a reduction in the elastic and inelastic interaction rates over time. In this paper, we introduce a momentum-dependent extension of the drag coefficient by using a first-order polynomial ansatz applied to collisional and radiative energy-loss coefficients. This approach dynamically links the drag force to the underlying energy-loss mechanisms, and goes beyond conventional models in which the drag coefficient is assumed to be momentum independent. Our framework enables a more detailed assessment of how momentum-dependent scattering rates influence collisional and radiative energy loss, and how they modify the nuclear modification factor of charm and bottom quarks.
It is important to clarify the phenomenological nature of this study. Our primary aim is to establish a baseline study that isolates the specific role of momentum dependence in the drag coefficient. To isolate this specific effect in a controlled setting, we employ several simplifying approximations. The drag coefficient is parametrized using a linear momentum expansion; the medium evolution is described by a Bjorken flow with a uniform transverse profile; phenomenological Fokker–Planck transport is applied; and an effective procedure is used to compare the results with D-meson data. These choices allow us to identify the genuine impact of the momentum-dependent terms, independently of model-specific hydrodynamic or hadronization uncertainties, providing a clear foundation for more complex and advanced studies of heavy-quark energy loss in future work.
@@
This paper is organized as follows: In the following section, we present a brief overview of the Fokker–Planck equation. We also discuss the space–time evolution of the system and specify the initial conditions employed in our analysis. In Section III, the drag and diffusion coefficients for collisional and radiative processes have been discussed. Our results are presented in Section IV, and section V is dedicated to summary and discussions.
@@ II System evolution
Heavy quarks reach thermodynamic equilibrium more slowly compared to lighter particles. Consequently, the interaction between heavy quarks and medium can be considered as interaction between equilibrium and non-equilibrium states. The equilibrium component, known as Quark-Gluon Plasma, is presumed to form at an initial temperature ($T_{i}$) and time ($\tau_{i}$) after the nuclear collision. The QGP expands and cools due to its high internal pressure, until transition to the hadronic phase at temperature $T_{c}$. Heavy quarks as the non-equilibrium component, exhibit Brownian motion within the thermal environment of the QGP Moore:2004tg ; Alam:1994sc . Therefore, Fokker-Planck equation, which is a simplified version of the Boltzmann equation for small momentum transformations Svetitsky:1987gq ; Rapp:2009my , provides a suitable framework for investigating the evolution of the heavy quark’s momentum distribution. According to this equation we have:
$\frac{\partial}{\partial t}f(p,t)=-\frac{\partial}{\partial p_{i}}\left[A_{i}(p)f(p,t)\right]+\frac{\partial^{2}}{\partial p_{i}\partial p_{j}}\left[D_{ij}(p)f(p,t)\right]$ (1)
Where $f(p,t)$ represents the distribution function of heavy quarks and $A_{i}(p)$ and $D_{ij}(p)$ are related to the drag and diffusion coefficients, respectively. These coefficients characterize the interaction of heavy quarks with the QGP and are generally influenced by both temperature and momentum. The i and j indices represent different spatial directions, which can be omitted as we are considering a spatially uniform quark-gluon plasma in transverse plane. In the next section, we will discuss these coefficients in detail.
@@
Where $T_{0}$ and $\tau_{0}$ represent the initial temperature and proper time, respectively. In this article, the initial proper time has been taken as $\tau_{0}=0.33~\text{fm}/c$ and the initial temperature has been set to $T_{0}=$ $403\text{\,}\mathrm{MeV}$. The parameter $\eta/s$ denotes the ratio of viscosity to entropy, which can be treated as a constant with a value of $1/4\pi$ Ruggieri:2013ova .
@@
We take $N_{f}=3$ as the number of active flavors present in the QGP, with the QCD cut-off parameter set at $\Lambda_{MS}=$ $80\text{\,}\mathrm{MeV}$.
The dominant influence of the medium’s temperature on plasma evolution is encoded in the temperature dependence of the strong coupling constant, as expressed in Eq. (3). However, other effects on the system dynamics arise through the temperature and momentum dependence of the drag and diffusion coefficients, which are discussed in the following section.
@@ III Drag and diffusion coefficients
To solve the Fokker-Planck equation and derive the distribution function of heavy quarks at the transition temperature ($T_{c}$), we require the drag and diffusion coefficients as a function of system parameters. It is important to note that the drag coefficient carries information about the dynamics of heavy quark interactions with the medium and is influenced by the properties of the thermal bath. Consequently, the key aspect of determining the temporal evolution of heavy quark distribution function lies in calculating the drag force acting on them, or equivalently, the rate of energy loss per unit distance traveled by the heavy quark in the medium. Thus, we can write:
@@
We consider both kinds of collisional and radiative energy loss in calculating the drag coefficient. Therefore, it can be determined through the following relation:
@@
The parameters $k_{1}$ and $k_{2}$ are typically treated as constants, which can be derived from the best fit to the experimental data. These constants incorporate the contributions from collision rates, degeneracy factors, and other relevant physical factors.
However, to precisely determine the drag coefficient, it is necessary to consider the impact of particles’ momenta on the rate of their collisions, which directly influences energy loss. As the kinetic energy and momentum of heavy quarks decrease during propagation in QGP, the frequency of their interactions with plasma constituents is expected to decrease as well. Therefore, the rates of both elastic and inelastic interactions depend on the momentum of the particles.
Consequently, we have formulated the drag coefficient as follows:
@@
The functions $k^{\prime}(p)$ and $k^{\prime\prime}(p)$ represent collisional and radiative energy loss coefficients, respectively. We assume these functions are comprised of a constant value and a modified term which is expressed as polynomial expansions of momentum, as follows:
$k^{\prime}(p)\>\approx\>k_{1}\>(1+\alpha\;p+O(P^{2})\>)$ (7)
$k^{\prime\prime}(p)\>\approx\>k_{2}\>(1+\beta\>p+O(P^{2})\>)$
We consider these expansions up to the first-order which include linear terms. Indeed we can ignore higher-order terms due to the lack of sufficient experimental data and their large uncertainties. By considering the energy loss coefficients as a function of $p$, the impact of particle’s momentum on the number of its interactions leading to collisional (elastic) and radiative (inelastic) energy losses can be examined separately. Setting $\alpha$ and $\beta$ to zero yields the original expression of the drag coefficient in the equation (5).
@@
All coefficients including $k_{1}$, $k_{2}$, $\alpha$, and $\beta$ should be determined by fitting to experimental data.
Once the drag coefficient has been determined, the diffusion coefficient can be obtained via Einstein’s relation, which is commonly applied in the weak-coupling or phenomenological context Diffusion.Srivastava:2016igg ; Diffusion.Das:2010tj ; Diffusion.Akamatsu:2008ge . According to this relation, we have:
$D(p)=TA(p)E$ (9)
Here, T is the temperature of the thermal bath and E is the heavy-quark energy. Therefore, by calculating the drag coefficient, we will have both the drag and diffusion coefficients.
It is important to emphasize that here the Einstein relation is employed purely as a phenomenological closure within the Fokker–Planck framework, rather than as a weak-coupling assumption derived microscopically from perturbative QCD. This approach is widely adopted in strongly-coupled heavy-quark transport studies Moore:2004tg ; vanHees:2004gq ; Das:2010tj and provides a consistent coarse-grained description of Brownian motion of heavy quarks in the QGP. The strong-coupling dynamics is effectively encoded in the momentum-dependent drag coefficient $A(p)$, whose functional form is constrained directly through comparison with experimental $R_{AA}$ data.
In order to calculate collisional and radiative energy loss, various methods have been proposed in recent studies. In our previous work Nezhad:2023mvb , we evaluated and compared three different approaches to calculate collisional energy loss and found that the approach proposed by Braaten and Thoma ( based on the Hard Thermal Loop (HTL) framework Braaten:1991we ) aligns better with recent experimental data. Therefore, we employ the same method for calculating collisional energy loss in our simulations. Radiative energy loss is calculated using the reaction operator formalism (DGLV) Gyulassy:2000fs ; Djordjevic:2003zk ; Wicks:2005gt and the generalized dead cone approach Saraswat:2015yll ; Saraswat:2017vuy .
Figure (1) presents the drag coefficient of charm quark for both cases of momentum-dependent coefficients and constant coefficients, at T=250MeVT=250\;MeV. As shown in the figure, incorporating the second-order expansion terms in the drag coefficient calculation leads to a significant increase in its value at higher momenta. Consequently, greater energy suppression is expected in these regions. By comparing the $R_{AA}$ results obtained from the above equation with those derived using constant energy loss coefficients (i.e., only the first term of the expansion), the degree of variation or improvement in the results can be quantified.
Figure 1: Drag coefficient as a function of $P_{T}$, calculated by constant rate (solid line) and momentum-dependent rate (dashed line) at T=250 MeV.
@@ IV Results
The nuclear suppression factor, $R_{AA}$, defined as the ratio of quenched to unquenched transverse momentum ($p_{T}$) distributions, quantifies the medium-induced energy loss of heavy quarks. It is influenced by several key processes, including the initial production of heavy quarks, the subsequent evolution of the quark-gluon plasma, and interactions of heavy quarks within the medium. In this work, we focus on the momentum-dependent drag coefficient and its effect on heavy-quark energy loss, formulating our study entirely at the heavy-quark level to establish a baseline quantification of drag-momentum dependence. A direct hadronization model, whether via fragmentation or coalescence, is not explicitly included. Instead, we employ an effective normalization procedure, a standard approach in baseline Fokker–Planck transport analyses Jamil:2010te ; Sheibani:2021ovo , which allows qualitative comparison with D-meson $R_{AA}$ while maintaining the focus on isolating the effect of the momentum-dependent drag.
The nuclear suppression factor can then be expressed as the ratio of the final heavy-quark distribution in heavy-ion collisions to the initial heavy-quark distribution in proton-proton collisions:
@@
The proportionality can be replaced with an equality by accounting for the number of binary collisions and the relevant degrees of degeneracy. This correction is implemented through an optimization procedure in our analysis.
In this section, we calculate the nuclear suppression factor of charm and bottom quarks in a Pb-Pb collision at a center-of-mass energy of 5.02 TeV and compare our findings with recent LHC experimental data collected in 2021 and 2022 ALICE:2020sjb ; ALICE:2021rxa ; ATLAS:2021xtw . The parton distribution functions are evolved using the Fokker–Planck equation, which we solve numerically on a Bjorken expanding medium, until the fireball reaches its freeze-out temperature of $T_{c}=155$ MeV. We calculate the drag coefficient using both cases of the first-order momentum expansion (8) and the constant rate approximation (5). This procedure allows us to compare the results with experimental data and investigate the impact of momentum dependence of energy loss coefficients on results.
The final outcomes are optimized to align with experimental data by adjusting free parameters such as $\alpha$, $\beta$, $k_{1}$ and $k_{2}$, by minimizing the unweighted Chi-squared value:
@@
$R_{AA}^{exp}$ and $R_{AA}^{th}$ represent experimental data and our theoretically calculated values for the suppression factor, respectively, while $\sigma$ indicates experimental error.
In our parameter optimization process, we use the Minuit package Minuit.James:1975dr , which is an effective and powerful tool that enables us to fit multiple free parameters and achieve high accuracy in minimizing the chi-squared value. This tool is essential for ensuring that the optimization is both precise and reliable, thereby enhancing the overall quality of the results.
Figure 2: Nuclear modification factor of charm quark in the (20–40%) centrality Pb-Pb collisions at $\sqrt{S_{NN}}=5.02$ TeV compared with ALICE 2021 data at forward rapidity ($2.5<y<4$) ALICE:2020sjb . The solid line represents the results obtained from the momentum-dependent coefficients, and dashed line corresponds to the drag coefficient with constant coefficients.
@@
Figures (2) and (3) demonstrate $R_{AA}$ results for charm quark, obtained by fitting theoretical predictions to the ALICE 2021 ALICE:2020sjb and ALICE 2022 ALICE:2021rxa datasets, respectively. A comparison of the two figures indicates that incorporating a momentum-dependent drag coefficient leads to better agreement with the experimental data, especially at high-momentum regions. This improvement is also supported by the $\chi^{2}/dof$ value which decreases from 0.3 to 0.2 for ALICE 2022 dataset and from 0.47 to 0.44 for ALICE 2021 dataset.
@@
The coefficients $\alpha$ and $\beta$ in Equation (8) for the c quark are found to be $\alpha=0.02$ and $\beta=0.05$. These values indicate that both collisional and radiative energy losses of particles increase effectively at higher momentum. Specifically, the impact of particle’s momentum on radiative energy loss becomes more significant and must be taken into account. We also obtain $k_{2}\approx 1.4\;k_{1}$ through the optimization process, which implies that the radiative energy loss experienced by charm quarks is consistently larger than collisional energy loss.
@@
Figure 4: Nuclear modification factor of bottom quark in the (40–60%) centrality Pb-Pb collisions at $\sqrt{S_{NN}}=5.02$ TeV (solid line) is compared with ATLAS collaboration data ATLAS:2021xtw and theoretical predictions from Shu-Qing Li et al. Li:2024wqq (dashed line).
Figure (4) shows our calculated $R_{AA}$ result for bottom quark obtained by fitting to the ATLAS 2022 data ATLAS:2021xtw . In this analysis, only the leading-order contribution to the drag coefficient was considered and the constant coefficients in Eq. (5) are found to satisfy approximately $k_{1}\approx 3\;k_{2}$. This hierarchy indicates a clear dominance of collisional over radiative energy loss for bottom quarks. Also, excellent agreement between theoretical predictions and experimental data is confirmed by our $\chi^{2}/dof$ value of 0.124, as illustrated in the figure (4). For further validation, we compare our calculated $R_{AA}$ values for bottom quark with Langevin-equation-based results reported in Ref. Li:2024wqq , also shown in Figure (4).
Due to the limited experimental data for the bottom quark, it is not feasible to accurately fit a model with too many free parameters. Therefore, investigating higher-order contributions to the drag coefficient is more challenging for bottom. Nevertheless, by repeatedly performing the fitting process, we can constrain the approximate bounds of the free parameters. By considering higher-order terms of drag coefficient, the parameters $\alpha$ and $\beta$ for this quark are found to be similar to those of charm quark. The key difference is that for the bottom quarks, even by considering higher order coefficients, we find $k_{1}$ greater than $k_{2}$, which indicates that for the bottom quark, collisional energy loss is the dominant mechanism even when higher-order terms are included. This is mainly due to the large mass of the bottom quark, which strongly suppresses gluon radiation Dokshitzer:2001zm .
Figure 5: Nuclear modification factor of bottom quark in the (40–60%) centrality Pb-Pb collisions at $\sqrt{S_{NN}}=5.02$ TeV compared with ATLAS 2022 data ATLAS:2021xtw .
Figure (5) presents the results derived from the momentum-dependent drag coefficients, for both upper and lower bounds of these coefficients, and compares them with the results based on constant coefficients. As shown in this figure, within the intermediate momentum range where experimental data are available, the drag coefficient with constant parameters provides satisfactory agreement with the experimental measurements. However, at higher momenta, the results obtained using momentum-dependent coefficients exhibit different behaviors, demonstrating a stronger suppression pattern. This observed discrepancy highlights the importance of future experimental data in the high-momentum regime to assess whether incorporating momentum dependence terms in the drag coefficient leads to more accurate theoretical predictions for bottom quarks.